OpenStudy (anonymous):

I GIVE MEDALS! How do I solve 100+23x=90(1.2^x) @Hero

OpenStudy (solomonzelman):

\(\large\color{slate}{ 100+23x=90(1.2^x) }\) this?!

OpenStudy (daniellelovee):

it has no solution

OpenStudy (solomonzelman):

no simple solution,... yes

OpenStudy (anonymous):

yes it has a solution

OpenStudy (anonymous):

Would the solution be the point of intersection if I graphed it on a coordinate plane?

OpenStudy (anonymous):

first solve 100+23x

OpenStudy (daniellelovee):

so far the farthest I got with this was 23x+100-9 times 1.2^x=0 which then equals 23x+1times1.2^x=0

OpenStudy (anonymous):

? @Daniellelovee

OpenStudy (solomonzelman):

just graphing both functions produces a lazy but simple to get result

OpenStudy (anonymous):

you're fast let her do it so she understands

OpenStudy (solomonzelman):

well, I did nothing besides putting it in a graphing calculator.

OpenStudy (freckles):

@georgia_catherine is the equation really: \[100+23x=90(1.2)^x\] like exactly how it is written?

OpenStudy (freckles):

Are we allowed to use any approximation methods to the intersections of the left hand expression and the right hand expression it will include knowing how to find the derivative or the method I have in mind like newton's method also there is no y

OpenStudy (anonymous):

im only in 8th grade

OpenStudy (solomonzelman):

when you say: \(\large\color{slate}{ 100+23x=90(1.2^x) }\) it's same as: \(\large\color{slate}{ y= 100+23x }\) \(\large\color{slate}{ y=90(1.2^x) }\)

OpenStudy (solomonzelman):

you can't solve it by a simple algebra: or even if you could it would hurt as hell.

OpenStudy (solomonzelman):

and substitution from that would get us our initial problem

OpenStudy (solomonzelman):

yes, I gave you the graph

OpenStudy (freckles):

are we suppose to use a method to approximate the solution?

OpenStudy (freckles):

like a table maybe?

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/src4elldzx

OpenStudy (solomonzelman):

I didn't zoom out to see that there is another solution-:( now I fixed that

Directrix (directrix):

@georgia_catherine Just wondering if the following is the original question you were asked to solve. s(x) = 100 + 23x m(x) = 90(1.2x) Complete the table of values. x 100 + 23x 90(1.2x) 0 1 2 3 4 5 Use the table to determine at approximately which point the number of cells will be the same for each organism. Graph the system of equations and show the point of intersection. Explain what the points graphed for each line represent. Explain how you can determine the solutio

Directrix (directrix):

From this URL: http://openstudy.com/updates/53d5fc48e4b081ba963c1884

Directrix (directrix):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) http://openstudy.com/updates/53e071f6e4b002c9d802bb9a @campbell_st worked on this before.